Wed Feb 08 2023

7mm issue

What would cause the 7MM numbers to Jump from (V4) to 3200-3300 whereas the 10mm stays consistent at around (V4) 2800. Fiber is good sliders are good. No issue with head or foc. lens. Replaced CPU same issue. Would the high voltage be an issue. Shorter pulse width is causing the problem, but adjustments are not working. Per Serv Tech.

Equipment: Candela VBeam classic

   RE: Surgical Laser Report Post

    This site uses cookies. By continuing to browse the site you are agreeing to our use of cookies. Please review our Privacy Policy for more details.
    I Agree